OpenStudy (anonymous):

Solve: 4x2 + 12x = 0 a) x = 0, -3 b) x = 3 c) x = 0, 3 d) x = -3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

\[4x^2+12x = 0\] You can divide both sides by 4 and get \[{4\over4} x^2 + {12\over4}x={0\over4}\]\[x^2+3x=0\]\[x(x+3)=0\] wich is true only if x is equal to 0 or x is equal to -3

OpenStudy (anonymous):

a is the answer